# Learning Quadrotor Control From Visual Features Using Differentiable Simulation

### v1

@@ IV-C Sim-to-real Transfer
Figure 7: Real-world trajectories showing stabilization from a manual throw.
For real-world experiments, we use a lightweight drone built with off-the-shelf components. Using hardware-in-the-loop (HITL) simulation, the pose of the physical quadrotor is monitored by a motion capture system, and the actor receives virtual observations, illustrated in Figure 1. The HITL experiments allow us to focus on testing the control performance without the need for a feature detection module. Note that even though the simulated observation model causes no sim-to-real gap, the measurement noise of the motion capture systems results in out-of-distribution observations.
Figure 6 illustrates the deployment of a vision-based control policy trained in differentiable simulation. The policy successfully navigates and stabilizes the quadrotor at the desired goal position. Figure 1 shows a second experiment where the quadrotor is thrown manually resulting in diverse and highly unstable initial states. However, as indicated by Figure 7, both state-based and vision-based control policies successfully perform the task in successive trials without failure.
@@ IV-D1 Policy Optimization Using Surrogate Gradients
Using a simpler dynamics model for the backward path trivially increases the number of simulation steps per second, as shown in Table I. However, using only a surrogate gradient might weaken the quality of the gradient and hence, decrease sample efficiency. Figure 8 shows two runs with BPTT over 200 iterations (100 environments) on the state-based control task. Both runs use the same initial parameters and random seeds. The only difference is the quadrotor model. It clearly shows that the simple model speeds up training but does not sacrifice sample efficiency.
IV-D2 State Representation Learning as Pretraining
@@
We hypothesize that state representation pretraining is particularly useful when combined with differentiable simulation because it not initializes the weights to provide a more meaningful internal representation but also improves the quality of the gradient. Possibly, the pretrained weights make the first-order approximation of the policy optimization objective more accurate in a larger vicinity of the current parameters.

### v2

@@ IV-C Sim-to-real Transfer
Figure 7: Real-world trajectories showing stabilization from a manual throw. Each line represents an individual trial.
For real-world experiments, we use a lightweight drone built with off-the-shelf components. Using hardware-in-the-loop (HITL) simulation, the pose of the physical quadrotor is monitored by a motion capture system, and the actor receives virtual observations, visualized in Figure 1. The HITL experiments allow us to focus on testing the control performance without the need for a feature detection module. Note that even though the simulated observation model causes no sim-to-real gap, the measurement noise of the motion capture systems results in out-of-distribution observations.
Figure 6 illustrates the deployment of a vision-based control policy trained in differentiable simulation. The policy successfully navigates and stabilizes the quadrotor at the desired goal position. Figure 1 shows a second experiment where the quadrotor is thrown manually resulting in diverse and highly unstable initial states. As indicated by Figure 7, both state-based and vision-based control policies successfully perform the task in successive trials without failure. The state-based policy stabilizes the quadrotor faster and with less oscillations around the desired position. In particular, in the third vision-based trial (denoted in red), the quadrotor came close to the floor before approaching the goal position. This behavior indicates that the vision-based performance could benefit from an improved observation noise model to reduce the sim-to-real gap.
@@ IV-D1 Policy Optimization Using Surrogate Gradients
Using a simpler dynamics model for the backward path directly increases the number of simulation steps per second, as shown in Table I. However, using only a surrogate gradient might weaken the quality of the gradient and hence, decrease sample efficiency. Figure 8 shows two runs with BPTT over 200 iterations (100 environments) on the state-based control task. Both runs use the same initial parameters and random seeds. The only difference is the quadrotor model. It clearly shows that the simple model speeds up training but does not sacrifice sample efficiency or performance.
IV-D2 State Representation Learning as Pretraining
@@
We hypothesize that state representation pretraining is particularly useful when combined with differentiable simulation because it not only initializes the weights to provide a more meaningful internal representation but also improves the quality of the gradient. Possibly, the pretrained weights make the first-order approximation of the policy optimization objective more accurate in a larger vicinity of the current parameters.